f(x)=1/2x^3-x
Find the zeros of the function algebraically

Answer by stanbon(75887)

f(x)=1/2x^3-x
Find the zeros of the function algebraically
Solve:
(1/2)x^3-x = 0
x((1/2)x^2-1) = 0
-----
x = 0 or (1/2)x^2 -1 = 0
-----
x = 0 or x^2 = 2
------
x = 0 or x = +/-sqrt(2)
--------------------
Cheers,
Stan H.
--------------